OpenStudy (anonymous):

writing two sentences how can 2x+2 to the power of 2 be corrected to make it equavilent to 2x+6

OpenStudy (anonymous):

@aum

OpenStudy (aum):

The question is not clear the way it appears. Can you take a screenshot of the question and post a link to it?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I agree with aum. Anyways, are you trying to do something like \[\Large (2x+2)^2 = 2x+6\]?

jimthompson5910 (jim_thompson5910):

Please post a screenshot of the problem. I have a feeling something is missing.

OpenStudy (anonymous):

i know im trying its hard

jimthompson5910 (jim_thompson5910):

its ok, I can wait for it to load

OpenStudy (anonymous):

do you have it????:/

jimthompson5910 (jim_thompson5910):

oooh ok, it was \[\Large 2x+2^2\], thx that helped

jimthompson5910 (jim_thompson5910):

\[\Large 2x + 2^2 = 2x + 4\] since \(\Large 2^2 = 4\), agreed?

jimthompson5910 (jim_thompson5910):

well you would write 2^2 as 2 squared. Or I guess you can keep that notation

jimthompson5910 (jim_thompson5910):

you want it equal to 2x+6, so how can you fix that equation?

jimthompson5910 (jim_thompson5910):

we have +4, but we want +6

jimthompson5910 (jim_thompson5910):

what can we do

OpenStudy (anonymous):

add two

jimthompson5910 (jim_thompson5910):

yes you add two to 4 to get 6

jimthompson5910 (jim_thompson5910):

so you have to add 2 to 2x+4 to get 2x+6

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thank you!!!!@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

you're welcome